OpenStudy (anonymous):

help 5 math questions and ill give medal and fan if corect Which related equation can be used to solve the equation? Column A Column B --- ABC 1. m + 5 = –7.3 --- ABC 2. m + 5 = 7.3 --- ABC 3. m + 7.3 = –5 --- ABC 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (anonymous):

wich one is wich the last 3 are the the choices

OpenStudy (phi):

I can't make sense of your table/question

OpenStudy (anonymous):

column a --abc 1. m+5=-7.3 --abc 2. m+5= 7.3 --abc 3. m+7.3=-5 --abc 4. m+7.3= 5

OpenStudy (anonymous):

colum b a is m=-7.3-5 b is m=5-7.3 c is m=7.3-5

OpenStudy (phi):

a in column b matches up with 1 in column a b with 4 c with 2 a matches up with 3

OpenStudy (anonymous):

thanks

OpenStudy (phi):

if you start with 1.) m + 5 = –7.3 add -5 to both sides m + 5 - 5 = -7.3 - 5 which simplifies to m = -7.3 - 5 which matches choice a in column b